OpenStudy (anonymous):

Determine which equation below has a solution of b = 3.

OpenStudy (anonymous):

A. 2b + 3b – 6 = –21 B. –8 = 3 + 4b – 8 – 5b C. –2 = 2 + 5b – 2 – 6b D. –4b + 3b – 6 = –4

OpenStudy (anonymous):

B 3+(4*3)-8-(5*3)=3+12-8-15=-8